OpenStudy (rarepepe):

how do you get the answer for this question?? 17 = –13 – 8x

OpenStudy (alexandervonhumboldt2):

add 13 to both sides: 17+13=-13-8x+13 30=-8x x=-(30/8)=simplify

OpenStudy (rarepepe):

thank you